A survey of 300 Senior High School learners shows that they read on average of 20 hours per week with a margin of error of 2.2 hours at 95% level of confidence.


21. What is the sample mean? A. 20 B. 300 C. 2.2 D. 0.05


22. Which of the following is the upper limit of the interval? A. 15 B. 17.8 C. 22.2 D. 1.96


23. What is the lower limit of the confidence interval? A. 17.8 B. 1.96 C. 22.2 D. 4.4


24. What is the length of the confidence interval? A. 17.8 B. 22.2 C. 4.4 D. 2.

1
Expert's answer
2022-06-12T15:07:20-0400

21. The sample mean is an average value found in a sample. In our sample it is 20 hours per week, so the answer is (a).


22. To find the upper limit of the confidence interval we have to add the margin of error to the mean:

"20+2.2=22.2,"

so the answer is (c).


23.To find the lower limit of the confidence interval we have to subtract the margin of error from the mean:

"20-2.2=17.8,"

so the answer is (a).


24. To find the length of the confidence interval, we have to double the margin of error:

"2.2\cdot 2=4.4,"

or to subtract the lower limit from the upper limit:

"22.2-17.8=4.4,"

so the answer is (c).


Answer: 21. (a) 22. (c) 23. (a) 24. (c)